OpenStudy (anonymous):

Infinite Geometric Series Sigma Problem will be posted below....I don't know how to find r...

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (2/3)(-3/4)^{n-1}\]

OpenStudy (rogue):

\[\sum_{n=1}^{\infty} \frac {2}{3} \times (\frac {-3}{4})^{n-1} = \sum_{n = 1}^{\infty} \frac {2}{3} \times \frac {-4}{3} \times = \frac {8}{9} \sum_{n=1}^{\infty} (\frac {-3}{4})^{n} = \frac {8}{9} \times \frac {\frac {-3}{4}}{1 - \frac {-3}{4}} = \frac {8}{9} \times \frac {-3}{7} = \frac {-8}{21}\]

OpenStudy (rogue):

= -8/21

OpenStudy (anonymous):

does \[a _{1}\] equal 2/3?

OpenStudy (anonymous):

because i thought you inserted the n=1 into the exponent of n-1? or am i wrong?

OpenStudy (anonymous):

The answer should be \[ \frac 8 {21}\] Note that \[ \sum_{n=1}^\infty a^{n-1} = \frac 1{1-a}, \quad |a| <1 \] Apply this with \[ a= -\frac 3 4 \] to your series and multiply what you get by \[ \frac 23\] to obtain \[ \frac 1 {28} \]

OpenStudy (anonymous):

so is the number before the exponent always r?

OpenStudy (anonymous):

Yes, but you should examine the first few terms. For example the following sum \[ \sum_{n=2}^\infty a^{n-1} = a + a^2 +a^3 +\cdots =\frac a{1-a}, \quad |a| <1\\ a \sum_{n=1}^\infty a^{n-1} = a \left (1+ a + a^2 +a^3 +\cdots\right) =a \left(\frac1{1-a}\right)=\frac a{1-a}, \quad |a| <1 \]

OpenStudy (rogue):

I have no idea what I typed, but yeah, the answer should be positive 8/21.